# Transfer Matrix Approach for Topological Edge States

### v3

@@ I Introduction
In this article, we introduce a new method to analyse one-dimensional topological insulators based on top of the well-established transfer matrix approach [17, 18, 19]. This transfer matrix method offers an approach complementary to the standard Bloch Hamiltonian method [20, 21]. As a preview, a system with $N$ sites per unit cell and $k$-nearest neighbour hoppings would admit a Bloch Hamiltonian of size $N\times N$, while the transfer matrix would have a size of $2k\times 2k$. In addition, while the Bloch Hamiltonian works best with periodic boundary conditions, the transfer matrix method can be naturally equipped with open or closed boundary conditions. A consequence of this is that the edge states appear naturally within the transfer matrix method, so that we may characterise the edge states more explicitly. We thus apply the transfer matrix method to inversion-symmetric models and obtain simple conditions for the appearance of their topologically protected edge states. This is done by uniquely parametrising the transfer matrix in terms of three real numbers using the Iwasawa decomposition of $\mathrm{SL}(2;\mathbb{R})$. This parametrisation reveals the submanifolds corresponding to symmetry-protected systems, and the topology of these submanifolds is intimately connected with the edge states of the system.
@@ II Transfer Matrix of the SSH model
When the chain is semi-infinite, this condition, along with the condition that $T^{k}(^{0}_{1})$ stays bounded in the chain, recovers the exact eigenmodes of the chain. This includes both the Bloch bands and the edge modes. It should be noted that the Bloch theorem no longer applies in the semi-infinite case as there is no translation symmetry in the system. To recover the exact eigenspectrum one needs to consider each mode as a superposition of counter-propagating waves in the infinite chain which fulfills the condition (6).
@@ III.2 Proof of the Zak Phase Bulk-Edge Correspondence
To connect between the transfer matrix picture and the Bloch Hamiltonian picture, let us first begin by analysing the Bloch eigenvector. Parity and time reversal [^3] The Hamiltonian considered in this paper has real entries, so it trivially satisfies time-reversal symmetry. This statemenet however is true for PT-symmetric Hamiltonians which are allowed to be complex. symmetry lets us write any Bloch eigenvector in the position basis as
@@ Appendix A Proof of Zak Phase-Eigenvector Phase relation
Consider a system with $N$ sites per unit cell. Without loss of generality, we may assume that $N$ is even. We label the amplitude in the $m^{\text{th}}$ site as $a_{m}$. In general, due to $PT$ symmetry, the Bloch eigenvector will satisfy
@@
$\left|K\right\rangle=\frac{\mathrm{e}\mkern 1.0mu^{\mathrm{i}\mkern 1.0mu\phi(K)/2}}{\sqrt{1+r_{2}^{2}(K)+\dots}}\begin{pmatrix}\mathrm{e}\mkern 1.0mu^{-\mathrm{i}\mkern 1.0mu\phi(K)/2}\\ r_{2}(K)\mathrm{e}\mkern 1.0mu^{-\mathrm{i}\mkern 1.0mu\phi_{2}(K)}\\ \vdots\\ r_{2}(K)\mathrm{e}\mkern 1.0mu^{\mathrm{i}\mkern 1.0mu\phi_{2}(K)}\\ \mathrm{e}\mkern 1.0mu^{\mathrm{i}\mkern 1.0mu\phi(K)/2}\end{pmatrix}.$ (31)
Note that the choice of reciprocal gauge $\mathrm{e}\mkern 1.0mu^{\mathrm{i}\mkern 1.0mu\phi(K)/2}$ is physical: It is required for the Zak phase to have an unambiguous interpretation of mean displacement. To prove this, we first fix the first site to have position $x_{1}=0$, so that the lattice position of the $m^{\text{th}}$ site is given by $x_{m}=m-1$. As a result, the Bloch eigenvector is related to the discrete fourier transform (DFT) by
@@
Importantly, the DFT of any state in the position basis is
[…]
As a result, for all $K$, we have that
$\langle m=1|K\rangle=\langle m=1|K,\text{DFT}\rangle=a_{1}=\text{const.}$. (34)
We may fix the global phase to 00, so that the Bloch eigenvector is given by (31). We thus see that allowing the phase of the first entry in $\left|K\right\rangle$ to vary is equivalent to choosing a different, translated coordinate system for each value of the system quasi momentum $K$.
We are now ripe to prove the assertion (12) in the main text. Defining $\left|K,\text{Sym}\right\rangle$ to be the RHS of (31) without the global phase factor, the action of the gauge transformation on the derivative is

### v4

@@ I Introduction
In this article, we introduce a new method to analyse one-dimensional topological insulators based on top of the well-established transfer matrix approach [17, 18, 19]. This transfer matrix method offers an approach complementary to the standard Bloch Hamiltonian method [20, 21]. As a preview, a system with $N$ sites per unit cell and $\ell$-nearest neighbour hoppings would admit a Bloch Hamiltonian of size $N\times N$, while the transfer matrix would have a size of $2\ell\times 2\ell$. In addition, while the Bloch Hamiltonian works best with periodic boundary conditions, the transfer matrix method can be naturally equipped with open or closed boundary conditions. A consequence of this is that the edge states appear naturally within the transfer matrix method, so that we may characterise the edge states more explicitly. We thus apply the transfer matrix method to inversion-symmetric models and obtain simple conditions for the appearance of their topologically protected edge states. This is done by uniquely parametrising the transfer matrix in terms of three real numbers using the Iwasawa decomposition of $\mathrm{SL}(2;\mathbb{R})$. This parametrisation reveals the submanifolds corresponding to symmetry-protected systems, and the topology of these submanifolds is intimately connected with the edge states of the system.
@@ II Transfer Matrix of the SSH model
For the remaining of this article, we will call any basis which satisfies the above condition a “position basis”. In the SSH model, this coincides exactly with the conventional meaning of the position basis, where each basis vector corresponds to a specific site on the lattice. When the chain is semi-infinite, this condition, along with the condition that $T^{k}(^{0}_{1})$ stays bounded in the chain, recovers the exact eigenmodes of the chain. This includes both the Bloch bands and the edge modes. It should be noted that the Bloch theorem no longer applies in the semi-infinite case as there is no translation symmetry in the system. To recover the exact eigenspectrum one needs to consider each mode as a superposition of counter-propagating waves in the infinite chain which fulfills the condition (6).
@@ III.2 Proof of the Zak Phase Bulk-Edge Correspondence
To connect between the transfer matrix picture and the Bloch Hamiltonian picture, let us first begin by analysing the Bloch eigenvector. Parity and time reversal [^3] The Hamiltonian considered in this paper has real entries, so it trivially satisfies time-reversal symmetry. This statemenet however is true for PT-symmetric Hamiltonians which are allowed to be complex. symmetry lets us write any Bloch eigenvector in the position basis (6) as
@@ Appendix A Proof of Zak Phase-Eigenvector Phase relation
Consider a system with $N$ sites per unit cell. Without loss of generality, we may assume that $N$ is even, since for a system with odd $N$, we can double the number of sites by merging 2 unit cells. We label the amplitude in the $m^{\text{th}}$ site as $a_{m}$. In general, due to $PT$ symmetry, the Bloch eigenvector will satisfy
@@
$\left|K\right\rangle=\frac{\mathrm{e}\mkern 1.0mu^{\mathrm{i}\mkern 1.0mu\phi(K)/2}}{\sqrt{1+r_{2}^{2}(K)+r_{3}^{2}(K)+\dots}}\begin{pmatrix}\mathrm{e}\mkern 1.0mu^{-\mathrm{i}\mkern 1.0mu\phi(K)/2}\\ r_{2}(K)\mathrm{e}\mkern 1.0mu^{-\mathrm{i}\mkern 1.0mu\phi_{2}(K)}\\ r_{3}(K)\mathrm{e}\mkern 1.0mu^{-\mathrm{i}\mkern 1.0mu\phi_{3}(K)}\\ \vdots\\ r_{3}(K)\mathrm{e}\mkern 1.0mu^{\mathrm{i}\mkern 1.0mu\phi_{3}(K)}\\ r_{2}(K)\mathrm{e}\mkern 1.0mu^{\mathrm{i}\mkern 1.0mu\phi_{2}(K)}\\ \mathrm{e}\mkern 1.0mu^{\mathrm{i}\mkern 1.0mu\phi(K)/2}\end{pmatrix}.$ (31)
Here, $r_{i}$ and $\phi_{i}$ are real-valued functions of $K$. Note that the choice of reciprocal gauge $\mathrm{e}\mkern 1.0mu^{\mathrm{i}\mkern 1.0mu\phi(K)/2}$ is physical: It is required for the Zak phase to have an unambiguous interpretation of mean displacement. To prove this, we first fix the first unit cell to have position 00. As a result, the Bloch eigenvector is related to the real-space Hamiltonian eigenstate as
$\left|K,\text{Real}\right\rangle=(1,e^{\mathrm{i}\mkern 1.0muK},e^{2\mathrm{i}\mkern 1.0muK},\cdots)\otimes\left|K\right\rangle$ (32)
As a result, for all $K$, we have that
$\langle 1|K,\text{Real}\rangle=a_{1}=\text{const.}$. (33)
Here, $\left|1\right\rangle$ is the position basis vector corresponding to the leftmost site, $K=2\pi m/L$ is the crystal momentum, $L$ is the number of unit cells, and $m$ is the unit cell number. We may fix the global phase to 00, so that the Bloch eigenvector is given by (31). We thus see that allowing the phase of the first entry in $\left|K\right\rangle$ to vary is equivalent to choosing a different, translated coordinate system for each value of the crystal momentum $K$.